# Magnetic levitation by rotation described by a new type of Levitron

### v1

@@ Abstract
Recently, the authors[1][2] have highlighted a novel magnetic levitation phenomenon utilizing two magnetically equivalent permanent neodymium magnets. We propose that this system represents an analog of a scaled-up version of a Levitron[3][4]. The main distinction between these systems lies in the fact that the radius $m_{/}\mu_{f}$ now is a function of the lateral displacement $\delta_{R}$, and the adjustment required for magnetic trapping is not determined by the rotation speed of the float. Using an effective model, we demonstrate that the observed levitation effect can be explained within the framework of [3][4], and the condition for the potential energy to be at a minimum at the trapping point is achieved with the existence of a constraint on the $\delta_{R}$ parameter.
Suggested keywords
I Introduction
@@
$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\ \frac{1}{2}\frac{\partial^{2}B_{z}}{\partial z^{2}}\left(\frac{1}{2B_{0}}\frac{(\frac{\partial B_{z}}{\partial z})^{2}}{\frac{\partial^{2}B_{z}}{\partial z^{2}}}-1\right)r^{2}+...)$ (1)
@@
In this work we will present a description of the phenomenon of rotational magnetic levitation discussed in Refs.[1][2] based on an effective model, considering an analogy with the model of a conventional Levitron described in Refs.[3][4]. Section II presents the equations that describe the motion of the floater, in section III based on considerations about the system’s potential energy, we establish the necessary conditions to explain the the magnetic trapping effect, observed in several available videos[7]. Section IV details our results, which are organized in the following three subsections. Finally in section V we present our conclusions.
II Neodymium floater dynamics
@@
where $U_{f}=-\vec{\mu}_{f}.\vec{B}_{R}$ In the absence of air resistance, or in the limit where $\xi^{R}_{f}=\xi^{T}_{f}=0$, Eq.(5) corresponds to equations (1) and (2) of Ref.[3] . In these equations, we will assume that the dynamics of the float is similar to that of a spinning top[3] composed of neodymium magnets.
@@ IV The decoupled case
FIG.(2a) shows the arrangement between the rotor (R) and the floater(f) in a scenario with a high degree of symmetry with respect to the z-axis, such that we can consider $\delta_{R}=0$ At this point, we highlight the differences between a conventional Levitron[3][4], and the so-called the ”dynamical Levitron”.
@@
and the above expression shows that for $\delta_{R}\to 0$, that condition (i) can be satisfied only if $13z_{f}/(z_{R}-z_{f})<$ 1.
@@
Therefore, even assuming that $13z_{f}/(z_{R}-z_{f})<$ 1, which is consistent with condition (i) , the above expression shows that the system is not able to satisfy both stability conditions simultaneously.
@@
In order to estimate the required value of $\delta_{R}$ to satisfy the stability conditions, FIG.(3) presents a contour plot of $U_{f}(z_{f})$ and $\partial^{2}_{z}U_{z}(z_{f})$, assuming $\delta_{R}=0.1mm$ for illustrative purposes. The y-axis of this figure represents $R_{f}$, while the x-axis represents $z_{f}(mm)$. To establish an adequate scale for this plot, the term $\partial^{2}_{z}U_{z}$ is multiplied by $10^{-5}$, where $U_{z}=U_{f}(z_{f})$.
In FIG(3a), it is evident that condition (i) is maintained, as indicated by the purple point. However, as shown in FIG.(3b), for the same region in the $(z_{f}(mm),R_{f})$ plane, condition (ii) is not simultaneously satisfied, and it can be verified that $\delta_{R}<\delta^{c}_{R}$.
@@
Figure 3: Contour plot obtained for Eqs.(9) and (12), for $\delta_{R}=0.1mm$. The contextualization of the curves behavior are described in the text.
In FIG.(3) we consider the choice $r_{f}=(14.2-z_{f})\,mm$, which is adequate to represent the results obtained for floaters formed by cylinders of height $h=5$, or spheres of diameter $D=(4)7$ as listed in the tables presented in Appendix A.
@@ V The coupled case
V.1 The minimum energy condition for the trapping
@@
In the videos available in[7], we demonstrate that this type of trapping is possible. Furthermore, based on direct measurements of the distance ($r_{f}$) between the rotor and the floater, in FIG.(4) we analyze the behavior of $U_{f}(z_{f})$ for $R_{f}=0$ for spherical, cylindrical, and cubic floaters as a function of the lateral displacement $\delta_{R}$. For the spherical floater, the cases $\delta_{R}\approx 1mm$, $1.5mm$ and $1.75mm$ are shown in FIGs. (4a)-(4c), respectively. For the cylindrical floater, the cases $\delta_{R}\approx 1mm$, $1.5mm$ are illustrated in FIGs. 4(d) and 4(e), while FIG. 4(f) presents the results for the cubic floater, which has the same dimensions as the cylinder, with $\delta_{R}=1.75mm$.
@@
The results presented in FIG.(4), obtained for different floaters and lateral displacements $\delta_{R}$, yield closely matching estimates for the critical value $\delta^{c}_{R}$. These findings suggest, that the minimum value $\delta^{c}_{R}$ required for trapping is of the order of $\sim O(0.3-0.4)mm$. Furthermore, they demonstrate that the magnetic alignment effect required for trapping is effectively achieved by a suitable choice of $\delta_{R}$, regardless of the geometry of the floater.
@@
Applying a straightforward dimensional analysis, we estimate that the critical value for the experimental apparatus employed in [1] would be of the order $\delta^{c}_{R}(19mm)\approx\frac{D}{h}\delta^{c}_{R}(5mm)\approx O(1)mm$ which agrees with our observations and highlights the sensitivity of $\delta^{c}_{R}$ to the geometry of the magnet.
In the following subsection, based on the series of measurements summarized in the tables of Appendix A, we will determine $\delta^{c}_{R}$ from Eq.(13) in order to compare with the results presented in this section.
@@
9.4 0.12 0.23
         cylinder 1.5mm 9.6 0.115 0.22
10 0.11 0.22
1.75mm 9.7 0.162 0.314
9.52 0.134 0.26
         sphere $(D=7mm)$ 1.5mm 9.74 0.125 0.243
9.9 0.12 0.24
9.76 0.134 0.26
         sphere $(D=4mm)$ 1,5mm 10.22 0.132 0.27
11.26 0.11 0.248
9.22 0.145 0.267
         sphere $(D=7mm)$ 1.75mm 9.60 0.129 0.248
10.03 0.113 0.23
Table 1: Experimental measurements arising from Tables (II-VII) for $r_{f}(mm)$ and $R_{f}$, with $\delta^{c}_{R}$ given by Eq.(15), for cylindrical and spherical floats with diameter and height of $D_{1}=7mm$, $D_{2}=4mm$ and $h=d=5mm$
V.2 The minimum energy condition for a dynamical trapping
In the previous section, we established the conditions for trapping a floater at rest, defined by $R_{f}\to 0$. We found that the minimum value of energy $U_{f}(z_{f})$, required to maintain this trapping, demands a minimum value assigned to the lateral displacement $\delta_{R}$, defined by $\delta^{c}_{R}\approx O(0.3-0.4)mm$.
Assuming that the floater is trapped, but now with angular velocity $\omega_{f}>0$, Eq.(13) yieds
@@ V.2 The minimum energy condition for a dynamical trapping
For $R_{f}>0$, in Eq.(14) the value assigned to $R_{f}\to(R_{f})_{min}$ corresponds to the situation where the floater has the minimum rotacional speed while being trapped, which consequently approaches the minimum value for $\delta_{R}$. At this point, the distance between the rotor and the floater $r_{f}$ is maximized approaching distance defined for $\delta^{c}_{R}$ when $R_{f}\approx 0$. Assuming the limit $r_{f}>>L$, it follows that
@@
Figure 5: Plot obtained for Eq.(9) assuming $(R_{f})_{min}$ for spherical float considering the data in Table I for $\delta_{R}=1.5mm$. The contextualization of the curves behavior are described in the text.
Eq.(15) is derived in the limit where the floater can be approximated as a point dipole. Corrections arising from finite size of the floater, parameterized by the multiplicative factor $(2+5L/r_{f})/(1+3L/r_{f})$, reduce the prediction of this equation by a factor $\lesssim O(0.1)mm$.
Based on this expression, it is possible to estimate $R_{f}(min)$ using the results presented in the previous section. The maximum distance observed for trapping of the floaters corresponds to $r_{f}\approx(10-11)mm$, which combined with $\delta^{c}_{R}\sim(0.3-0.4)mm$ implies
@@
Based on above expression, the rotational speed of the floater relative to the rotor, where $r_{f}\approx O(10)mm$, should satisfy $\omega_{f}\lesssim 0.17\omega_{R}$. The prediction $\omega_{f}\lesssim 0.17\omega_{R}$ is corroborated by the results reported in Tables (II-V) for $\omega_{R}$ and $\omega_{f}$ listed in Appendix A. A synthesis of these measurements, highlighting the results for cylindrical and spherical floaters, is provided in Table I.
@@
Finally, in Tables (VI-VII) present results for cases where $R_{f}>R_{f}(min)$ In these case, the $U_{f}(z_{f})$ curves which describe the trapping, are characterized by $\delta^{c}_{R}<dR_{f}/5$, corresponding to the region between the blue and red dash-dotted curves shown in FIG.(4), for a given rotor configuration. The observed increase in $R_{f}$ for example, compared to the data in Table III for distances of the same order arises from differences in the rotor configuration, specifically the lateral displacement employed.
In the upper left corner of FIGs.(4a)-(4b), it can be observed that the rotor frequency required for trapping the same floater for the cases where $\delta_{R}\approx 1mm$ and $\delta_{R}\approx 1.5mm$, are $\omega_{R}=12,330\,rpm$ and $\omega_{R}=8,946\,rpm$,respectively. As discussed previously, in this type of device, angular momentum is transferred from the rotor to the floater. For example, in the case where $\delta_{R}\approx 1mm$ Table VII shows that the respective increase in $R_{f}$ results from the greater transfer of angular momentum from the rotor.
For the results presented in Table VI, where the lateral displacement is given by $\delta_{R}\approx 1.75\text{ mm}$, a comparison can be made with the data from Table II. Since the energy available for trapping the floats is proportional to $\delta_{R}$, for this rotor configuration we have more energy available, so the trapping for $r_{f}\approx(9.6-9.7)\text{ mm}$ occurs at a lower rotation speed $\omega_{R}\approx^{{}^{\!\!\!\!\!VI}}O(11,100)\text{ rpm}$ when compared to the trapping observed for $\delta_{R}\approx 1.5\text{ mm}$, where $\omega_{R}\approx^{{}^{\!\!\!\!\!II}}O(12,500)\text{ rpm}$; the indices $(VI,II)$ refer to the respective tables.
Therefore, from the discussion raised in the above paragraphs, it is possible to verify that the increase in lateral displacement $\delta_{R}$, when we compare different configurations $\delta^{A}_{R}>\delta^{B}_{R}$, is in principle, responsible for the magnetic trapping of the float observed at rotor speeds of the order of $O(10,000)rpm$ for the cases reported in FIGs.(4b),(4c),(4e) and (4f). In addition, by $\omega^{{}^{\!VI}}_{R}<\omega^{{}^{\!II}}_{R}$ discussed previously. Thus, an important question arises ”Is there an upper limit for $\delta_{R}=\delta^{max}_{R}$, beyond which the effect of magnetic trapping ceases to occur?”.
@@
V.3 The condition on the maximum allowed value of lateral displacement

### v3

@@ Abstract
Recently, a novel magnetic levitation phenomenon involving two magnetically equivalent neodymium permanent magnets has been reported. In this work, we propose that this system functions as a scaled-up analog of the Levitron. The key distinction is that the ratio $m_{/}\mu_{f}$ becomes a function of the lateral displacement $\delta_{R}$, and magnetic trapping no longer depends on the rotational speed of the levitating body as in a conventional Levitron. Furthermore, we demonstrate that stable trapping occurs when a specific constraint on the $\delta_{R}$ parameter is satisfied, ensuring that the potential energy reaches a minimum at the equilibrium point.
Keywords: Magnetism, magnetic levitation, levitron, magnetic trap
1 Introduction
@@
[…]
@@
In Refs.[1][2], the authors established the fundamental physical principles underlying the magnetic levitation observed in conventional Levitron systems. In such systems, stable levitation requires a precise tuning of the radius $m_{/}\mu_{f}$, which in turn determines the allowable range of rotation frequencies for the spinning top, ensuring that the system reaches a potential minimum at the trapping point.
In this work, we extend that framework to the case of a dynamical Levitron, demonstrating that the effective radius $m_{/}\mu_{f}$ becomes modulated by $\delta_{R}$. This modulation provides a natural explanation for the magnetic levitation by rotation described in Refs.[4][6]. Furthermore, it allows us to identify the necessary conditions under which the system evolves toward a minimum energy configuration, thereby enabling stable magnetic trapping. Section II presents the equations that describe the motion of the floater, in section III based on considerations about the system’s potential energy, we establish the necessary conditions to explain the magnetic trapping effect. Sections IV and V details our results, which are organized into five subsections . Finally in section VI we present our conclusions.
2 Neodymium floater dynamics
@@
where $U_{f}=-\vec{\mu}_{f}.\vec{B}_{R}$ . In the absence of air resistance, or in the limit where $\xi^{R}_{f}=\xi^{T}_{f}=0$, Eq.(5) corresponds to equations (1) and (2) of Ref.[1] . In these equations, we will assume that the dynamics of the float is similar to that of a spinning top[1] composed of neodymium magnets.
@@ 4 The decoupled case
FIG.(2a) shows the arrangement between the rotor (R) and the floater(f) in a scenario with a high degree of symmetry with respect to the z-axis, such that we can consider $\delta_{R}=0$ . At this point, we highlight the differences between a conventional Levitron[1][2], and the so-called the ”dynamical Levitron”.
@@
and the above expression shows that for $\delta_{R}\to 0$, that condition (i) can be satisfied only if $3z_{f}/(z_{R}-z_{f})<1$.
@@
Therefore, even assuming that $3z_{f}/(z_{R}-z_{f})<1$, which is consistent with condition (i) , the above expression shows that the system is not able to satisfy both stability conditions simultaneously.
@@
In order to estimate the required value of $\delta_{R}$ to satisfy the stability conditions, FIG.(3) presents a contour plot of $U_{f}(z_{f})$ and $\partial^{2}_{z_{f}}U_{f}(z_{f})$, assuming $\delta_{R}=0.1mm$ for illustrative purposes. The y-axis of this figure represents $R_{f}$, while the x-axis represents $z_{f}(mm)$. To establish an adequate scale for this plot, the term $\partial^{2}_{z_{f}}U_{f}$ is multiplied by $10^{-5}$.
In FIG(3a), it is evident that condition (i) is maintained, as indicated by the purple point. This region, depicted in blue, represents an area of stability, while the instability region is delimited in red. Therefore, as shown in FIG.(3b), for the same region in the $(z_{f}(mm),R_{f})$ plane, condition (ii) is not simultaneously satisfied, and it can be verified that $\delta_{R}<\delta^{c}_{R}$.
@@
Figure 3: Contour plot obtained from Eq.(9) for $\delta_{R}=0.1mm$. The contextualization of the curves behavior are described in the text.
In FIG.(3) we consider the choice $r_{f}=(14.2-z_{f})\,mm$, which is adequate to represent the results obtained for floaters formed by cylinders of height $h=5$, or spheres of diameter $D=(4)7$ as listed in the tables presented in Appendix B.
@@ 5 The coupled case
5.1 The minimum energy condition for the trapping: Floater at rest
@@
5.2 The experimental determination of $\delta^{c}_{R}$ for an floater at rest
The experimental setup employed in this work is thoroughly detailed in Appendix A. As described therein, our study involved the recording of the phenomenon and precise measurements of the rotor and floater rpm, utilizing a digital tachometer and high-resolution camera operating at 960 frames per second (fps). A rigorous analysis of these videos was subsequently performed using the Tracker software. All video recordings of the experiments conducted in this study, which document the experimental observations of trapping, are archived and made publicly available in the repository referenced in [7]. Among the diverse conditions under which trapping was observed and recorded, we particularly highlight the possibility of trapping occurring even when the floater remains at rest. The quantitative results extracted from this and all other related analyses are systematically compiled in tables presented in Appendix B.
Based on direct measurements of the distance ($r_{f}$) between the rotor and the floater, in FIG.(4) we analyze the behavior of $U_{f}(z_{f})$ for $R_{f}=0$ for spherical, cylindrical, and cubic floaters as a function of the lateral displacement $\delta_{R}$. For the spherical floater, the cases $\delta_{R}\approx 1mm$, $1.5mm$ and $1.75mm$ are shown in FIGs. (4a)-(4c), respectively. For the cylindrical floater, the cases $\delta_{R}\approx 1mm$, $1.5mm$ are illustrated in FIGs. 4(d) and 4(e), while FIG. 4(f) presents the results for the cubic floater, which has the same dimensions as the cylinder, with $\delta_{R}=1.75mm$.
@@ 5.2 The experimental determination of $\delta^{c}_{R}$ for an floater at rest
The experimental results presented in FIG.(4), obtained for different floaters and lateral displacements $\delta_{R}$, yield closely matching estimates for the critical value $\delta^{c}_{R}$. These findings suggest, that the minimum value $\delta^{c}_{R}$ required for trapping is of the order of $\sim O(0.3-0.4)mm$. Furthermore, they demonstrate that the magnetic alignment effect required for trapping is effectively achieved by a suitable choice of $\delta_{R}$, regardless of the geometry of the floater.
The results found for $\delta_{R}$ are in agreement with the Eq.(13), which establishes that in the limit where $R_{F}\to 0$, that the trapping condition (ii) can only be satisfied for $\delta_{R}>0$. However, it is important to emphasize that in this limit Eq.(13) is not able to predict a specific value for $\delta^{c}_{R}$, with $\delta^{c}_{R}$ being defined by the minimum energy of the float required to maintain trapping, where the energy of the float $E_{f}$ corresponds to $U(0)_{grav}$.
@@
The results reported in Figs.(4a-4f) are from floaters with diverse geometries. Consequently, for a more precise determination for $\delta^{c}_{R}$ it would be necessary to characterize the non-uniformity of magnetic moments inherent for example to cylindrical and cubic floaters that would go beyond the scope of this work, which aims to establish the basic conditions necessary to explain the observed magnetic trapping.
@@
Applying a straightforward dimensional analysis, we estimate that the critical value for the experimental apparatus employed in [6] would be of the order $\delta^{c}_{R}(19mm)\approx\frac{D}{h}\delta^{c}_{R}(5mm)\approx O(1)mm$ which agrees with our observations and is also in accordance with recent results reported in [8] highlights the sensitivity of $\delta^{c}_{R}$ to the geometry of the magnet.
Table 1: Experimental measurements arising from Tables (B1-B6) for $r_{f}(mm)$ and $R_{f}$, with $\delta^{c}_{R}$ given by Eq.(15), for cylindrical and spherical floats with diameter and height of $D_{1}=7mm$, $D_{2}=4mm$ and $h=d=5mm$ .
@@
[…]
$\pm 0.1$ $1±\pm 0$.0011 $2±\pm 0$.0022
[…]
In the following subsection, based on the series of measurements summarized in the tables of Appendix B, we will determine $\delta^{c}_{R}$ from Eq.(13) assuming $R_{F}>0$, which allows in this case to extract $\delta^{c}_{R}$ directly from Eq.(13) in order to compare with the results presented in this section.
5.3 The minimum energy condition for a dynamical trapping: The minimum rotational energy of the floater
In the previous section, we established the conditions for trapping a floater at rest, defined by $R_{f}\to 0$. We found that the minimum value of energy $U_{f}(z_{f})$, required to maintain this trapping, demands a minimum value assigned to the lateral displacement $\delta_{R}$, defined by $\delta^{c}_{R}\approx O(0.3-0.4)mm$. Now, assuming that floater has the minimum rotational energy we will, based on experimental measurements for $R_{f}$, verify the agreement of the predictions of Eq.(13) with the results previously obtained.
Assuming that the floater is trapped, but now with angular velocity $\omega_{f}>0$, Eq.(13) yields
@@ 5.3 The minimum energy condition for a dynamical trapping: The minimum rotational energy of the floater
For $R_{f}>0$, in Eq.(14) the value assigned to $R_{f}\to(R_{f})_{min}$ corresponds to the situation where the floater has the minimum rotational speed while being trapped, which consequently approaches the minimum value for $\delta_{R}$. At this point, the distance between the rotor and the floater $r_{f}$ is maximized approaching distance defined for $\delta^{c}_{R}$ when $R_{f}\approx 0$. Assuming the limit $r_{f}>>L$, it follows that
@@
5.4 The experimental determination of $\delta^{c}_{R}$ for an dynamical trapping
Based on the experimental measurements of $r_{f}$ and $R_{f}$ extracted from Tables B1B6B1-B6, we directly determine $\delta^{c}_{R}$ using Eq.(15) . In Table 1, the calculated $\delta^{c}_{R}$ values range from $\delta^{c}_{R}=(0.23-0.31)$mm , which are in good agreement with previous findings. It is important to note that Eq.(15) is derived under the point-dipole approximation for the floater. Finite-size corrections, parameterized by the multiplicative factor $(2+5L/r_{f})/(1+3L/r_{f})$, reduce the prediction of this equation by a factor $\lesssim O(0.1)mm$.
Having established a reasonable agreement between the $\delta^{c}_{R}$ values obtained directly from Eq.(15) and previous results, we can now proceed to characterize the necessary conditions for the minimum rotational energy of the floater, denoted as $R_{f}(min)$. Utilizing the results from the previous section, specifically $\delta^{c}_{R}\sim(0.3-0.4)$mm , and the observation that the maximum trapping distance for the floaters corresponds to $r_{f}\approx(10-11)$mm, we then obtain
Figure 5: Plot obtained for Eq.(9) assuming $(R_{f})_{min}$ for spherical float considering the data in Table 1 for $\delta_{R}=1.5mm$. The contextualization of the curves behavior are described in the text.
@@ 5.4 The experimental determination of $\delta^{c}_{R}$ for an dynamical trapping
Based on above expression, the rotational speed of the floater relative to the rotor, where $r_{f}\approx O(10)mm$, should satisfy $\omega_{f}\lesssim 0.17\omega_{R}$. The prediction $\omega_{f}\lesssim 0.17\omega_{R}$ is corroborated by the results reported in Tables (B1-B4) for $\omega_{R}$ and $\omega_{f}$ listed in Appendix B.
@@
The $\delta^{c}_{R}$ values obtained are in good agreement with those observed for spherical floaters, as depicted in Figs. 4a-4c. However, it is crucial to consider that dynamic trapping introduces the effect of dynamic drag. This drag can lead to a deviation between the predictions of Eq.(15) (derived under ideal conditions) and the experimentally observed $\delta^{c}_{R}$ values. A direct comparison reveals a deviation of approximately $\sim O(20\%)$, which is directly attributable to drag effects. Notably, this magnitude of deviation is consistent with that obtained in previous experiments carried out by the authors[1].
Finally, in Tables (B5-B6) present results for cases where $R_{f}>R_{f}(min)$ . In these case, the $U_{f}(z_{f})$ curves which describe the trapping, are characterized by $\delta^{c}_{R}<dR_{f}/5$, corresponding to the region between the blue and red dash-dotted curves shown in FIG.(4), for a given rotor configuration. The observed increase in $R_{f}$ for example, compared to the data in Table B2 for distances of the same order arises from differences in the rotor configuration, specifically the lateral displacement employed.
In the upper left corner of FIGs.(4a)-(4b), it can be observed that the rotor frequency required for trapping the same floater for the cases where $\delta_{R}\approx 1mm$ and $\delta_{R}\approx 1.5mm$, are $\omega_{R}=12,330\,rpm$ and $\omega_{R}=8,946\,rpm$,respectively. As discussed previously, in this type of device, angular momentum is transferred from the rotor to the floater. For example, in the case where $\delta_{R}\approx 1mm$ Table B6 shows that the respective increase in $R_{f}$ results from the greater transfer of angular momentum from the rotor.
For the results presented in Table B5, where the lateral displacement is given by $\delta_{R}\approx 1.75$mm, a comparison can be made with the data from Table B1. Since the energy available for trapping the floats is proportional to $\delta_{R}$, for this rotor configuration we have more energy available, so the trapping for $r_{f}\approx(9.6-9.7)$mm occurs at a lower rotation speed $\omega_{R}\approx^{{}^{\!\!\!\!\!B5}}O(11,100)$ rpm when compared to the trapping observed for $\delta_{R}\approx 1.5$mm, where $\omega_{R}\approx^{{}^{\!\!\!\!\!B1}}O(12,500)$ rpm; the indices $(B5,B1)$ refer to the respective tables.
Therefore, from the discussion raised in the above paragraphs, it is possible to verify that the increase in lateral displacement $\delta_{R}$, when we compare different configurations $\delta^{A}_{R}>\delta^{B}_{R}$, is in principle, responsible for the magnetic trapping of the float observed at rotor speeds of the order of $O(10,000)rpm$ for the cases reported in FIGs.(4b),(4c),(4e) and (4f). In addition, by $\omega^{{}^{\!B5}}_{R}<\omega^{{}^{\!B1}}_{R}$ discussed previously. Thus, an important question arises ”Is there an upper limit for $\delta_{R}=\delta^{max}_{R}$, beyond which the effect of magnetic trapping ceases to occur?”.
@@
5.5 The condition on the maximum allowed value of lateral displacement